# On the Theoretical Limitations of Embedding-Based Retrieval

### v1

@@ Abstract
Vector embeddings have been tasked with an ever-increasing set of retrieval tasks over the years, with a nascent rise in using them for reasoning, instruction-following, coding, and more. These new benchmarks push embeddings to work for any query and any notion of relevance that could be given. While prior works have pointed out theoretical limitations of vector embeddings, there is a common assumption that these difficulties are exclusively due to unrealistic queries, and those that are not can be overcome with better training data and larger models. In this work, we demonstrate that we may encounter these theoretical limitations in realistic settings with extremely simple queries. We connect known results in learning theory, showing that the number of top-$k$ subsets of documents capable of being returned as the result of some query is limited by the dimension of the embedding. We empirically show that this holds true even if we restrict to $k=2$, and directly optimize on the test set with free parameterized embeddings. We then create a realistic dataset called LIMIT that stress tests models based on these theoretical results, and observe that even state-of-the-art models fail on this dataset despite the simple nature of the task. Our work shows the limits of embedding models under the existing single vector paradigm and calls for future research to develop methods that can resolve this fundamental limitation.
\@testdef
undefined
@@ 1 Introduction
00footnotetext: Work done during internship at GDM.00footnotetext: Data and code are available at https://github.com/google-deepmind/limit
@@
Over the last two decades, information retrieval (IR) has moved from models dominated by sparse techniques (such as BM25 (Robertson et al., 1995)) to those that use neural language models (LM) as their backbones (Lee et al., 2019; Craswell et al., 2020; Izacard et al., 2021; Wang et al., 2022). These neural models are predominantly used in a single vector capacity, where they output a single embedding representing the entire input (also known as dense retrieval). These embedding models are capable of generalizing to new retrieval datasets and have been tasked with solving increasingly complicated retrieval problems (Thakur et al., 2021; Enevoldsen et al., 2025; Lee et al., 2025).
@@
In recent years this has been pushed even further with the rise of instruction-following retrieval benchmarks, where models are asked to represent any relevance definition for any query (Weller et al., 2025a, b; Song et al., 2025; Xiao et al., 2024; Su et al., 2024). For example, the QUEST dataset (Malaviya et al., 2023) uses logical operators to combine different concepts, studying the difficulty of retrieval for complex queries (e.g., “Moths or Insects or Arthropods of Guadeloupe”). On the other hand, datasets like BRIGHT (Su et al., 2024) explore the challenges stemming from different definitions of relevance by defining relevance in ways that require reasoning. One subtask includes reasoning over a given Leetcode problem (the query) to find other Leetcode problems that share a sub-task (e.g. others problems using dynamic programming). Although models cannot solve these benchmarks yet, the community has proposed these problems in order to push the boundaries of what dense retrievers are capable of—which is now implicitly every task that could be defined.
@@
Rather than proposing empirical benchmarks to gauge what embedding models can achieve, we seek to understand at a more fundamental level what the limitations are. Since embedding models use vector representations in geometric space, there exists well-studied fields of mathematical research (Papadimitriou and Sipser, 1982) that could be used to analyze these representations.
@@
Our work aims to bridge this gap, connecting known theoretical results in geometric algebra with modern advances in neural information retrieval. We draw upon research in communication complexity theory to provide a lower bound on the embedding dimension needed to represent a given combination of relevant documents and queries. Specifically, we show that for a given embedding dimension $d$ there exists top-$k$ combinations of documents that cannot be returned—no matter the query—highlighting a theoretical and fundamental limit to embedding models.
@@
Figure 1: A depiction of the LIMIT dataset creation process, based on theoretical limitations. We test all combinations of relevance for $N$ documents (i.e. in the figure, all combinations of relevance for three documents with two relevant documents per query) and instantiate it using a simple mapping. Despite this simplicity, SoTA MTEB models perform poorly, scoring less than 20 recall@100.
@@
To show that this theoretical limit is true for any retrieval model or training dataset, we test a setting where the vectors themselves are directly optimized with the test data. This allows us to empirically show how the embedding dimension enables the solving of retrieval tasks. We find there exists a crucial point for each embedding dimension ($d$) where the number of documents is too large for the embedding dimension to encode all combinations. We then gather these crucial points for a variety of $d$ and show that this relationship can be modeled empirically with a polynomial function.
@@
We also go one step further and construct a realistic but simple dataset based on these theoretical limitations (called LIMIT). Despite the simplicity of the task (e.g., who likes Apples? and Jon likes Apples, ...), we find it is very difficult for even state-of-the-art embedding models (Lee et al., 2025; Zhang et al., 2025) on MTEB (Enevoldsen et al., 2025) due to the theoretical underpinnings, and impossible [^1] At least with current optimization techniques for retrieval. for models with small embedding dimensions.
@@
These results imply interesting findings for the community: on one hand we see neural embedding models becoming immensely successful. However, academic benchmarks test only a small amount of the queries that could be issued (and these queries are often overfitted to), hiding these limitations. Our work shows that as the tasks given to embedding models require returning ever-increasing combinations of top-$k$ relevant documents (e.g., through instructions connecting previously unrelated documents with logical operators), we will reach a limit of combinations they cannot represent.
Thus, the community should be aware of these limitations, both when designing evaluations (as LIMIT shows) and by choosing alternative retrieval approaches – such as cross-encoders or multi-vector models – when attempting to create models that can handle the full range of instruction-based queries, i.e. any query and relevance definition.
@@ 2.1 Neural Embedding Models
There has been immense progress on embedding models in recent years (Lee et al., 2019; Craswell et al., 2020; BehnamGhader et al., 2024), moving from simple web search (text-only) to advanced instruction-following and multi-modal representations. These models generally followed advances in language models, such as pre-trained LMs (Hoffmann et al., 2022), multi-modal LMs (Li et al., 2024; Team, 2024), and advances in instruction-following (Zhou et al., 2023; Ouyang et al., 2022). Some of the prominent examples in retrieval include CoPali (Faysse et al., 2024) and DSE (Ma et al., 2024) which focus on multimodal embeddings, Instructor (Su et al., 2022) and FollowIR (Weller et al., 2024a) for instruction following, and GritLM (Muennighoff et al., 2024) and Gemini Embeddings (Lee et al., 2025) for pre-trained LMs turned embedders.
@@ 2.2 Empirical tasks pushing the limits of dense retrieval
Retrieval models have been pushed beyond their initial use cases to handle a broad variety of areas. Notable works include efforts to represent a wide group of domains (Thakur et al., 2021; Lee et al., 2024), a diverse set of instructions (Weller et al., 2024a; Zhou et al., 2024; Oh et al., 2024), and to handle reasoning over the queries (Xiao et al., 2024; Su et al., 2024). This has pushed the focus of embedding models from basic keyword matching to embeddings that can represent the full semantic meaning of language. As such, it is more common than ever to connect what were previously unrelated documents into the top-$k$ relevant set, [^2] You can imagine an easy way to connect any two documents merely by using logical operators, i.e. X and Y. increasing the number of combinations that models must be able to represent. This has motivated our interest in understanding the limits of what embeddings can represent, as current work expects it to handle every task.
Previous work has explored empirically the limits of models: Reimers and Gurevych (2020) showed that smaller dimension embedding models have more false positives, especially with larger-scale corpora. Ormazabal et al. (2019) showed the empirical limitations of models in the cross-lingual setting and Yin and Shen (2018) showed how embedding dimensions relate to the bias-variance tradeoff. In contrast, our work provides a theoretical connection between the embedding dimension and the sign-rank of the query relevance (qrel) matrix, while also showing empirical limitations.
@@ 2.3 Theoretical Limits of Vectors in Geometric Space
Understanding and finding nearest neighbors in semantic space has a long history in mathematics research, with early work such as the Voronoi diagram being studied as far back as 1644 and formalized in 1908 (Voronoi, 1908). The order-k version of the Voronoi diagram (i.e. the Voronoi diagram depicting the set of closest k points) is an obvious analog to information retrieval and has been studied for many years (Clarkson, 1988). However, proofs placing a bound on the count of the number of regions in the order-k Voronoi problem are notoriously different to bound tightly and do not provide much practical insight for IR (Bohler et al., 2015; Lee, 1982; Chen et al., 2023).
We approach this problem from another angle by proving that the set of the constraints implied by the top-$k$ retrieval problem can be formalized to show that it places a lower bound on the dimensionality of the embedding needed to represent it. We then show that this dimensionality can be much larger than the dimensionality of embedding models for practical IR problems. This approach relies on previous work in the communication complexity theory community to place bounds using the sign-rank of a matrix. Due to the difficulty of of computing the sign-rank, we rely on known previous work that has already proven the sign-rank of known matrices (Hatami et al., 2022; Alon et al., 2014; Chierichetti et al., 2017; Chattopadhyay and Mande, 2018; Hatami and Hatami, 2024). Our results also provide a proof of a method that can place a lower bound on the sign rank through what we call free embeddings in §4 (i.e. if it can be solved, then the dimension $d$ is $\leq$ to the sign rank).

### v2

@@ Abstract
Vector embeddings have been tasked with an ever-increasing set of retrieval tasks over the years, with a nascent rise in using them for reasoning, instruction-following, coding, and more. These new benchmarks push embeddings to work for any query and any notion of relevance that could be given. While prior works have pointed out theoretical limitations of vector embeddings, there is a common assumption that these difficulties are exclusively due to unrealistic queries, and those that are not can be overcome with better training data and larger models. In this work, we demonstrate that we may encounter these theoretical limitations in realistic settings with extremely simple queries. We connect known results in learning theory, showing that the number of top-$k$ subsets of documents capable of being returned as the result of some query is limited by the dimension of the embedding. We empirically show that this holds true even if we directly optimize on the test set with free parameterized embeddings. We then create a realistic dataset called LIMIT that stress tests embedding models based on these theoretical results, and observe that even state-of-the-art models fail on this dataset despite the simple nature of the task. Our work shows the limits of embedding models under the existing single vector paradigm and calls for future research to develop new techniques that can resolve this fundamental limitation.
@@ 1 Introduction
Over the last two decades, information retrieval (IR) has moved from models dominated by sparse techniques (such as BM25 Robertson et al. (1995)) to those that use neural language models (LM) as their backbones (Lee et al., 2019; Craswell et al., 2020; Izacard et al., 2021; Wang et al., 2022). These neural models are predominantly used in a single vector capacity, where they output a single embedding representing the entire input (also known as dense retrieval). These embedding models are capable of generalizing to new retrieval datasets and have been tasked with solving increasingly complicated retrieval problems (Thakur et al., 2021; Enevoldsen et al., 2025; Lee et al., 2025).
@@
In recent years this has been pushed even further with the rise of instruction-following retrieval benchmarks, where models are asked to represent any relevance definition for any query (Weller et al., 2025a; c; Song et al., 2025; Xiao et al., 2024; Su et al., 2024). For example, the QUEST dataset (Malaviya et al., 2023) uses logical operators to combine different concepts, studying the difficulty of retrieval for complex queries (e.g., “Moths or Insects or Arthropods of Guadeloupe”). On the other hand, datasets such as BRIGHT (Su et al., 2024) explore the challenges arising from different definitions of relevance by defining relevance in ways that require reasoning. One subtask includes reasoning over a given Leetcode problem (the query) to find other Leetcode problems that share a subtask (e.g. others problems using dynamic programming). Although models cannot solve these benchmarks yet, the community has proposed these problems in order to push the boundaries of what dense retrievers are capable of—which is now implicitly every task that could be defined.
@@
Rather than proposing empirical benchmarks to gauge what embedding models can achieve, we seek to understand at a more fundamental level what the limitations are. Since embedding models use vector representations in geometric space, there exist well-studied fields of mathematical research (Papadimitriou & Sipser, 1982) that could be used to analyze these representations.
@@
Our work aims to bridge this gap, connecting known theoretical results in linear algebra with modern advancements in neural information retrieval. We draw upon research in high-dimensional geometry to provide a lower bound on the embedding dimension needed to represent a given combination of relevant documents and queries. Specifically, we show that for a given embedding dimension $d$ there exist top-$k$ combinations of documents that cannot be returned—no matter the query—highlighting a theoretical and fundamental limit to embedding models.
@@
Figure 1: A depiction of the LIMIT dataset creation process, based on theoretical limitations. We test all combinations of relevance for $N$ documents (i.e. in the figure, all combinations of relevance for three documents with two relevant documents per query) and instantiate it using a simple mapping.
@@
To show that this theoretical limit is true for any retrieval model or training dataset, we test a setting where the vectors themselves are directly optimized with the test data. This allows us to empirically show how the embedding dimension enables the solving of retrieval tasks. We find that there exists a crucial point for each embedding dimension ($d$) where the number of documents is too large for the embedding dimension to encode all combinations. We then gather these crucial points for a variety of $d$ and show that this relationship can be modeled empirically with a polynomial function.
@@
We also go one step further and construct a realistic but simple dataset based on these theoretical limitations (called LIMIT). [^1] Data and code are available at https://github.com/google-deepmind/limit Despite the simplicity of the task (e.g., who likes Apples? and Jon likes Apples, ...), we find it is very difficult for even state-of-the-art embedding models (Lee et al., 2025; Zhang et al., 2025) on MTEB (Enevoldsen et al., 2025), and practically impossible for models with small embedding dimensions using standard optimization techniques.
@@
These results imply interesting findings for the community: on one hand we see neural embedding models becoming immensely successful. However, academic benchmarks test only a small amount of the queries that could be issued (and these queries are often overfitted to), hiding these limitations. Our work shows that as the tasks given to embedding models require returning ever-increasing combinations of top-$k$ relevant documents (e.g., through instructions connecting previously unrelated documents with logical operators), we will reach a limit of combinations they can represent.
Thus, the community should be aware of these limitations, both when creating evals and also by using alternate architectures—such as cross-encoders / multi-vector / more expressive similarity functions —when trying to handle the full range of instruction queries, i.e. any query and relevance definition.
@@ 2.1 Neural Embedding Models
There has been immense progress on embedding models in recent years (Lee et al., 2019; Craswell et al., 2020; BehnamGhader et al., 2024), moving from simple web search (text-only) to advanced instruction-following and multi-modal representations. These models generally followed advancements in language models, such as pre-trained LMs (Hoffmann et al., 2022), multi-modal LMs (Li et al., 2024; Team, 2024), and advancements in instruction-following (Zhou et al., 2023; Ouyang et al., 2022). Some of the prominent examples in retrieval include CoPali (Faysse et al., 2024) and DSE (Ma et al., 2024) which focus on multimodal embeddings, Instructor (Su et al., 2022) and FollowIR (Weller et al., 2024a) for instruction following, and GritLM (Muennighoff et al., 2024) and Gemini Embeddings (Lee et al., 2025) for pre-trained LMs turned embedders.
@@ 2.2 Empirical tasks pushing the limits of dense retrieval
Retrieval models have been pushed beyond their initial use cases to handle a broad variety of areas. Notable works include efforts to represent a wide group of domains (Thakur et al., 2021; Lee et al., 2024), a diverse set of instructions (Weller et al., 2024a; Zhou et al., 2024; Oh et al., 2024; Weller et al., 2025b), and to handle reasoning over the queries (Xiao et al., 2024; Su et al., 2024). This has pushed the focus of embedding models from basic keyword matching to embeddings that can represent the full semantic meaning of language. As such, it is more common than ever to connect what were previously unrelated documents into the top-$k$ relevant set, [^2] You can imagine an easy way to connect any two documents merely by using logical operators, i.e. X and Y. increasing the number of combinations that models must be able to represent. This has motivated our interest in understanding the limits of what embeddings can represent, as current work expects it to handle every task.
Previous work has explored empirically the limits of models: Reimers & Gurevych (2020) showed that smaller dimension embedding models have more false positives, especially with larger-scale corpora. Ormazabal et al. (2019) showed the empirical limitations of models in the cross-lingual setting and Yin & Shen (2018) showed how embedding dimensions relate to the bias-variance tradeoff. In contrast, our work provides a theoretical connection between the embedding dimension and the top-k sets it can retrieve, while also showing empirical limitations.
@@ 2.3 Theoretical Limits of Vectors in Geometric Space
Understanding and finding nearest neighbors in semantic space has a long history in mathematics research, with early work such as the Voronoi diagram being studied as far back as 1644 and formalized in 1908 (Voronoi, 1908). The order-$k$ version of the Voronoi diagram (i.e. the Voronoi diagram partitioning the space into regions based on their closest $k$ points) is obviously connected to information retrieval and has been studied for many years (Clarkson, 1988). The number of such regions is equal to the number of unique retrieval sets of size $k$, however this quantity is notoriously difficult to bound tightly (Bohler et al., 2015; Lee, 1982; Chen et al., 2023).
We approach this problem from a different angle, asking not how many $k$-subsets a given configuration realizes, but rather what embedding dimension is necessary to realize all $k$-subsets with a guaranteed score margin. By applying a classical sphere-packing volume argument (Vershynin, 2018; Conway et al., 1999), we obtain a lower bound on the embedding dimension in terms of $n$, $k$, and the margin $\gamma$. Our result is conceptually related to the Johnson–Lindenstrauss lemma (Johnson et al., 1984), which gives a sufficient dimension to preserve pairwise distances among $n$ points; in contrast, our bound gives a necessary dimension to realize all retrieval sets with a margin. The role of the margin in controlling the complexity of realizable configurations also parallels classical results in statistical learning theory, including the fat-shattering dimension (Kearns & Schapire, 1994) and margin-based generalization bounds for linear classifiers (Bartlett, 2002; Vapnik, 1998), where larger margins similarly constrain the capacity of the hypothesis class.